# A Unified View of Optimal Kernel Hypothesis Testing

### v1

@@ 5.2 Uniform power against alternatives separated in HSIC metric
$\max_{k\in{\mathcal{K}}}\,\max_{\ell\in{\mathcal{L}}}\,\mathrm{HSIC}_{k,\ell}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \max\biggl{\{}\sqrt{\frac{\max\{\log(1/\alpha),\log(1/\beta),\log(|{\mathcal{K}}||{\mathcal{L}}|)\}}{N}},\ \frac{r}{N}\biggr{\}}.$ (35)
@@ 5.3 Uniform power against alternatives separated in L2 metric
For the independence kernel tests, we present uniform separation rates in terms of the $L^{2}$-norm of the difference between the joint $p_{xy}$ and the product of the marginals $p_{x}\otimes p_{y}$, with Sobolev regularity assumption on $p_{xy}-p_{x}\otimes p_{y}$. Translation-invariant kernels are used and their bandwiths varied. Minimax optimal rates can be attained by using the optimal kernel bandwidth which depends on the unknown Sobolev smoothness (i.e., cannot be implementable). Adaptivity over the unknown Sobolev smoothness $s$ can be achieved using aggregation with multiple testing (Section 2) over the kernel bandwidths independently of $s$.
@@
The HSIC test, with optimal kernel bandwidths depending on the unknown Sobolev smoothness $s$, controls the type II error by $\beta$ for alternatives satisfying (Albert et al., 2022, Corollary 2 with theoretical quantiles, and Kim et al., 2022, Proposition 8.7 with permuted quantiles, see also Section A.7)
$\|p_{xy}-p_{x}\otimes p_{y}\|_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N}\right)^{2s/(4s+d)}$ (36)
@@
To avoid the dependence on the unknown Sobolev smoothness $s$, we resort to aggregation (multiple testing) over a collection of pairs of bandwidths, of size $(\log N)^{2}$, independent of $s$. The resulting test achieves the minimax optimal rate up to an iterated logarithmic term (Albert et al., 2022, Corollary 3 with theoretical quantiles, and Schrab et al., 2022b, Theorem 3 with estimated quantiles, see also Section A.9)
$\|p_{xy}-p_{x}\otimes p_{y}\|_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N/\log(\log(N))}\right)^{2s/(4s+d)}.$ (37)
@@
The test based on the efficient HSIC estimator with design $|{\mathcal{D}}|$ and fixed kernels is powerful provided that (Section A.8 and Schrab et al., 2022b, Theorem 1)
$\|p_{xy}-p_{x}\otimes p_{y}\|_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{|{\mathcal{D}}|/N}\right)^{2s/(4s+d)}.$ (38)
When $|{\mathcal{D}}|\asymp N^{2}$, the rate matches the standard minimax optimal rate. As $|{\mathcal{D}}|$ decreases, it deteriorates until it no longer converges for linear tests with $|{\mathcal{D}}|\asymp N$. This uniform separation rate holds assuming that optimal kernel bandwidths are used, as these depend on the unknown Sobolev smoothness $s$, this test cannot be implemented in practice. To overcome this issue, one can use aggregation over the bandwidths (multiple testing), losing the unwanted dependence on $s$. This results in the same power guarantee up to an iterated logarithmic term (Section A.9 and Schrab et al., 2022b, Theorem 2)
$\|p_{xy}-p_{x}\otimes p_{y}\|_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{\big{(}|{\mathcal{D}}|/N\big{)}\big{/}\log\!\big{(}\!\log(|{\mathcal{D}}|/N)\big{)}}\right)^{2s/(4s+d)}.$ (39)
@@
The differentially private dpHSIC test (Kim and Schrab, 2023, Algorithm 1 and Appendix B.5) achieves different uniform separation rates depending on the value of $\xi=\varepsilon+\log\!\big{(}{1}/({1-\delta})\big{)}$ compared to rates in $N$ (Section A.11 and Kim and Schrab, 2023, Theorem 14). This creates three different privacy regimes. In the low privacy regime with $\xi\gtrsim N^{-(2s-d/2)/(4s+d)}$, dpHSIC achieves the non-DP minimax rate
$\|p_{xy}-p_{x}\otimes p_{y}\|_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ N^{-2s/(4s+d)},$ (40)
@@
$\|p_{xy}-p_{x}\otimes p_{y}\|_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ (N^{3/2}\xi)^{-s/(2s+d)}.$ (41)
@@
$\|p_{xy}-p_{x}\otimes p_{y}\|_{L^{2}}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \left(N\xi\right)^{-2s/(2s+d)}.$ (42)
As in the two-sample case, these rates logarithmically depend on $\alpha$ and we believe that a logarithmic dependence in $\beta$ can also be obtained (Section A.11). Deriving matching lower bounds for $L^{2}$ separation under differential privacy constraint is an open problem, left for future work.
@@ 6 Goodness-of-fit testing
Finally, we consider a third framework: non-parametric goodness-of-fit testing. In Section 6.1, we define the goodness-of-fit testing framework and its associated wild-bootstrap which allows to compute a quantile and, hence, also to construct a well-calibrated test. In Section 6.2 and Section 6.3, we provide power guarantees in terms of KSD and $L^{2}$ Sobolev uniform separation rates, respectively, under standard, efficiency, privacy and robustness constraints. We refer the reader to Schrab (2025b, Section 2.3) for a detailed introduction to the Kernel Stein Discrepancy (KSD).
@@ 6.1 Framework, bootstrap and level
Given access to a model distribution $P$ and to i.i.d. samples $X_{1},\dots,X_{N}$ from a distribution $Q$, the aim is to test whether the two distributions are equal, that is, ${\mathcal{H}}_{0}\colon P=Q$, or not, i.e., ${\mathcal{H}}_{1}\colon P\neq Q$. This goodness-of-fit problem is sometimes referred to as one-sample testing. As in the two-sample case, the general setting of Section 1 is covered by having ${\mathcal{P}}$ as the space of all pairs of distributions, ${\mathcal{P}}_{0}$ as $\{(P,Q)\in{\mathcal{P}}:P=Q\}$, and ${\mathcal{P}}_{1}$ as $\{(P,Q)\in{\mathcal{P}}:P\neq Q\}$. We use the notation ${\mathbb{X}}_{N}\coloneqq(X_{1},\dots,X_{N})$.
@@
The type of access to the model distribution $P$ can differ depending on the testing setting. If a simulator allowing to sample from $P$ is available, then the goodness-of-fit problem essentially reduces to the two-sample problem where as many samples as desired can be requested from the model distribution. Being able to sample from the model, essentially simulating the null hypothesis, is closely related to the notion of parametric bootstrap (Stute et al., 1993). In other cases, it might not be possible to have access to a simulator, but we can have some knowledge about the model distribution itself. For example, it can sometimes be possible to compute the kernel expectations under the model in closed form, in which case a test can be constructed using the one-sample MMD plug-in estimator $\mathrm{MMD}_{k}^{2}(P,\widehat{Q})$ (with the MMD as expressed in Schrab, 2025b, Equation 6) which is equal to
@@
Consider the two-sample testing problem where we are given i.i.d. samples $X_{1},\dots,X_{m}$ from a distribution $P$, and i.i.d. samples $Y_{1},\dots,Y_{n}$ from a distribution $Q$, all independent from each other, and we are interested in testing whether $P=Q$. As the model distribution, we can consider the empirical distribution $\widehat{P}$ (uniform distribution on the samples). In this setting, we are then able to compute expectations under the model (empirical) distribution. Running a goodness-of-fit test specialised for this setting can then be used to solve the original two-sample problem. For example, we can use the one-sample MMD estimator of Equation 43 (corresponding to the KSD estimator with the simple Stein operator of Schrab, 2025b, Equation 58) with the model empirical distribution $\widehat{P}$. It clear that computing the expectations with respect to $\widehat{P}$ in this estimator then leads to the usual MMD estimator.
@@
To construct a KSD test following the procedure of Section 1.2, we need a method to simulate the null when using a KSD one-sample second-order statistic (see Schrab, 2025b, Section 3). For this, we rely on the wild bootstrapped statistics (see Section 1.2 for details) which can be computed as ${|{\mathcal{D}}|}^{-1}\sum_{(i,j)\in{\mathcal{D}}}\varepsilon_{i}\varepsilon_{j}h_{P}(X_{i},X_{j})$ where $\varepsilon_{1},\dots,\varepsilon_{n}$ are realisations of i.i.d. Rademacher variables, and $h_{P}$ is the Stein kernel as defined in Schrab (2025b, Equation 43). The KSD test resulting from the test construction of Section 1.2 with the wild bootstrap is then guaranteed to control the type I error asymptotically (Chwialkowski et al., 2014, Theorem 1). The wild bootstrapped statistics can all be computed efficiently as outlined in Section 1.2.
@@
The KSD test is known to be consistent achieving pointwise power (Chwialkowski et al., 2016, Proposition 3.2, and Liu et al., 2016, Proposition 4.2), that is, for any fixed alternative, the power of the KSD test converges to 1 asymptotically. In Sections 6.2 and 6.3, we present asymptotic power guarantees in a setting in which alternatives depend on the sample size and shrink towards the null as it increases.
@@
The permutation-based privatisation and robustisation procedures of Kim and Schrab (2023, Algorithm 1) and Schrab and Kim (2025, Algorithm 1) only work in frameworks where testing the null corresponds exactly to testing exchangeability (see Section 1). As previously mentioned, the two-sample and independence testing frameworks satisfy this property. However, this is not the case for goodness-of-fit testing. Indeed, since we have access to only one sample (and to the model itself), we are not able to permute samples across distributions as in the two-sample case, and permuting the points within one sample has no effect. So, the goodness-of-fit problem cannot be framed as testing for exchangeability. This is the reason why there are no permutation method in this setting, and why the type I error control only holds asymptotically. This also explains why the procedures of Kim and Schrab (2023, Algorithm 1) and Schrab and Kim (2025, Algorithm 1) cannot be used to construct private and robust KSD tests: a task which remains an open problem.
@@
The choice of base kernel and its bandwidth leads to drastically different Stein kernels, the role of the bandwidth is even more crucial in the goodness-of-fit setting as it affects the derivates of the base kernel which appear in the expression of the Stein kernel. Hence, this choice greatly impacts the power of the KSD test. This kernel selection problem can be solved with the aggregation method (Section 2) or with kernel pooling (Schrab, 2025b, Section 4). We recommend using these two adaptive procedures, leading to the KSDAgg test (Schrab et al., 2022a, Algorithm 1) and the normalised KSDFuse test.
@@ 6.2 Uniform power against alternatives separated in KSD metric
The KSD test is guaranteed to have power at least $1-\beta$ against all alternatives separated as (Section A.1)
$\mathrm{KSD}_{k}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sqrt{\frac{\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta)\big{\}}}{N}}$ (44)
which is of order $N^{-1/2}$. Using any unnormalised mean kernel pooling (Schrab, 2025b, Section 4) leads to (Section A.4)
$\underset{k\in{\mathcal{K}}}{\mathrm{mean}}\leavevmode\nobreak\ \mathrm{KSD}_{k}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sqrt{\frac{\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta)\big{\}}}{N}},$ (45)
while using unnormalised fuse/max kernel pooling leads to (Section A.4)
$\max_{k\in{\mathcal{K}}}\,\mathrm{KSD}_{k}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sqrt{\frac{\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta),\,\log(|{\mathcal{K}}|)\big{\}}}{N}},$ (46)
which, for a collection of kernels of size $|{\mathcal{K}}|=\log N$ (e.g., Schrab et al., 2022a, Theorem 3.5.ii), corresponds to a rate of order $(N/\log\log N)^{-1/2}$. Here, and throughout this section, the fusing parameter $\nu$ is assumed to be greater than $N$ and $\log(|{\mathcal{K}}|)$. In the following, we present results for fuse/max kernel pooling, while mean pooling results can be derived in a similar way to Equation 45 achieving the same rate as for fixed kernel.
@@
The efficient goodness-of-fit test, based on the block KSD B-statistic (Schrab, 2025b, Equation 70) consisting of $B$ blocks, with fixed kernel $k$, controls the type II error by $\beta$ provided that (Section A.5)
$\mathrm{KSD}_{k}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sqrt{\frac{B\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta)\big{\}}}{N}}.$ (47)
If only one block is used (i.e., $B=1$ corresponding to the complete U-statistic), then the same rate $N^{-1/2}$ as in Equation 44 is achieved. As the number of blocks $B$ increases, the rate $(B/N)^{1/2}$ gets slower until it no longer converges to zero for $B\asymp N$. The unnormalised fuse/max pooled block KSD statistic over a collection ${\mathcal{K}}$ of kernels (often of size $|{\mathcal{K}}|=\log(|{\mathcal{D}}|/N)\approx\log(N/B)$ as in Schrab et al. (2022b, Theorem 2.ii) with $|{\mathcal{D}}|=B\lfloor N/B\rfloor^{2}\asymp N^{2}/B$) achieves the same uniform separation rate with an additional $\sqrt{\log|{\mathcal{K}}|}$ term, that is (Section A.4)
$\max_{k\in{\mathcal{K}}}\,\mathrm{KSD}_{k}\leavevmode\nobreak\ \gtrsim\leavevmode\nobreak\ \sqrt{\frac{B\max\!\big{\{}\!\log(1/\alpha),\,\log(1/\beta),\,\log(|{\mathcal{K}}|)\big{\}}}{N}}.$ (48)

### v2

@@ 5.2 Uniform power against alternatives separated in HSIC metric
$\max_{k\in{\mathcal{K}}}\,\max_{\ell\in{\mathcal{L}}}\,\mathrm{HSIC}_{k,\ell}\penalty 10000\ \gtrsim\penalty 10000\ \max\biggl\{\sqrt{\frac{\max\{\log(1/\alpha),\log(1/\beta),\log(|{\mathcal{K}}||{\mathcal{L}}|)\}}{N}},\ \frac{r}{N}\biggr\}.$ (35)
@@ 5.3 Uniform power against alternatives separated in L2 metric
For the independence kernel tests, we present uniform separation rates in terms of the $L^{2}$-norm of the difference between the joint $p_{xy}$ and the product of the marginals $p_{x}\otimes p_{y}$, with Sobolev regularity assumption on $p_{xy}-p_{x}\otimes p_{y}$. Translation-invariant kernels are used and their bandwiths varied. Minimax optimal rates can be attained by using the optimal kernel bandwidth which depends on the unknown Sobolev smoothness (i.e., cannot be implementable). Adaptivity over the unknown Sobolev smoothness $s$ can be achieved using aggregation with multiple testing (Section˜2) over the kernel bandwidths independently of $s$.
@@
The HSIC test, with optimal kernel bandwidths depending on the unknown Sobolev smoothness $s$, controls the type II error by $\beta$ for alternatives satisfying (Albert et al., 2022, Corollary 2 with theoretical quantiles, and Kim et al., 2022, Proposition 8.7 with permuted quantiles, see also Section˜A.7)
$\|p_{xy}-p_{x}\otimes p_{y}\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N}\right)^{2s/(4s+d)}$ (36)
@@
To avoid the dependence on the unknown Sobolev smoothness $s$, we resort to aggregation (multiple testing) over a collection of pairs of bandwidths, of size $(\log N)^{2}$, independent of $s$. The resulting test achieves the minimax optimal rate up to an iterated logarithmic term (Albert et al., 2022, Corollary 3 with theoretical quantiles, and Schrab et al., 2022b, Theorem 3 with estimated quantiles, see also Section˜A.9)
$\|p_{xy}-p_{x}\otimes p_{y}\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{N/\log(\log(N))}\right)^{2s/(4s+d)}.$ (37)
@@
The test based on the efficient HSIC estimator with design $|{\mathcal{D}}|$ and fixed kernels is powerful provided that (Section˜A.8 and Schrab et al., 2022b, Theorem 1)
$\|p_{xy}-p_{x}\otimes p_{y}\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{|{\mathcal{D}}|/N}\right)^{2s/(4s+d)}.$ (38)
When $|{\mathcal{D}}|\asymp N^{2}$, the rate matches the standard minimax optimal rate. As $|{\mathcal{D}}|$ decreases, it deteriorates until it no longer converges for linear tests with $|{\mathcal{D}}|\asymp N$. This uniform separation rate holds assuming that optimal kernel bandwidths are used, as these depend on the unknown Sobolev smoothness $s$, this test cannot be implemented in practice. To overcome this issue, one can use aggregation over the bandwidths (multiple testing), losing the unwanted dependence on $s$. This results in the same power guarantee up to an iterated logarithmic term (Section˜A.9 and Schrab et al., 2022b, Theorem 2)
$\|p_{xy}-p_{x}\otimes p_{y}\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ \!\left(\frac{\log(1/\alpha)\log(1/\beta)}{\big(|{\mathcal{D}}|/N\big)\big/\log\!\big(\!\log(|{\mathcal{D}}|/N)\big)}\right)^{2s/(4s+d)}.$ (39)
@@
The differentially private dpHSIC test (Kim and Schrab, 2023, Algorithm 1 and Appendix B.5) achieves different uniform separation rates depending on the value of $\xi=\varepsilon+\log\!\big({1}/({1-\delta})\big)$ compared to rates in $N$ (Section˜A.11 and Kim and Schrab, 2023, Theorem 14). This creates three different privacy regimes. In the low privacy regime with $\xi\gtrsim N^{-(2s-d/2)/(4s+d)}$, dpHSIC achieves the non-DP minimax rate
$\|p_{xy}-p_{x}\otimes p_{y}\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ N^{-2s/(4s+d)},$ (40)
@@
$\|p_{xy}-p_{x}\otimes p_{y}\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ (N^{3/2}\xi)^{-s/(2s+d)}.$ (41)
@@
$\|p_{xy}-p_{x}\otimes p_{y}\|_{L^{2}}\penalty 10000\ \gtrsim\penalty 10000\ \left(N\xi\right)^{-2s/(2s+d)}.$ (42)
As in the two-sample case, these rates logarithmically depend on $\alpha$ and we believe that a logarithmic dependence in $\beta$ can also be obtained (Section˜A.11). Deriving matching lower bounds for $L^{2}$ separation under differential privacy constraint is an open problem, left for future work.
@@ 6 Goodness-of-fit testing
Finally, we consider a third framework: non-parametric goodness-of-fit testing. In Section˜6.1, we define the goodness-of-fit testing framework and its associated wild-bootstrap which allows to compute a quantile and, hence, also to construct a well-calibrated test. In Section˜6.2 and Section˜6.3, we provide power guarantees in terms of KSD and $L^{2}$ Sobolev uniform separation rates, respectively, under standard, efficiency, privacy and robustness constraints. We refer the reader to Schrab (2025b, Section 2.3) for a detailed introduction to the Kernel Stein Discrepancy (KSD).
@@ 6.1 Framework, bootstrap and level
Given access to a model distribution $P$ and to i.i.d. samples $X_{1},\dots,X_{N}$ from a distribution $Q$, the aim is to test whether the two distributions are equal, that is, ${\mathcal{H}}_{0}\colon P=Q$, or not, i.e., ${\mathcal{H}}_{1}\colon P\neq Q$. This goodness-of-fit problem is sometimes referred to as one-sample testing. As in the two-sample case, the general setting of Section˜1 is covered by having $\mathcal{P}$} as the space of all pairs of distributions, ${\mathcal{P}}_{0}$ as $\{(P,Q)\in{\mathcal{P}}:P=Q\}$, and ${\mathcal{P}}_{1}$ as $\{(P,Q)\in{\mathcal{P}}:P\neq Q\}$. We use the notation ${\mathbb{X}}_{N}\coloneqq(X_{1},\dots,X_{N})$.
@@
[…]
@@
Consider the two-sample testing problem where we are given i.i.d. samples $X_{1},\dots,X_{m}$ from a distribution $P$, and i.i.d. samples $Y_{1},\dots,Y_{n}$ from a distribution $Q$, all independent from each other, and we are interested in testing whether $P=Q$. As the model distribution, we can consider the empirical distribution $\widehat{P}$ (uniform distribution on the samples). In this setting, we are then able to compute expectations under the model (empirical) distribution. Running a goodness-of-fit test specialised for this setting can then be used to solve the original two-sample problem. For example, we can use the one-sample MMD estimator of Equation˜43 (corresponding to the KSD estimator with the simple Stein operator of Schrab, 2025b, Equation 58) with the model empirical distribution $\widehat{P}$. It clear that computing the expectations with respect to $\widehat{P}$ in this estimator then leads to the usual MMD estimator.
@@
To construct a KSD test following the procedure of Section˜1.2, we need a method to simulate the null when using a KSD one-sample second-order statistic (see Schrab, 2025b, Section 3). For this, we rely on the wild bootstrapped statistics (see Section˜1.2 for details) which can be computed as ${|{\mathcal{D}}|}^{-1}\sum_{(i,j)\in{\mathcal{D}}}\varepsilon_{i}\varepsilon_{j}h_{P}(X_{i},X_{j})$ where $\varepsilon_{1},\dots,\varepsilon_{n}$ are realisations of i.i.d. Rademacher variables, and $h_{P}$ is the Stein kernel as defined in Schrab (2025b, Equation 43). The KSD test resulting from the test construction of Section˜1.2 with the wild bootstrap is then guaranteed to control the type I error asymptotically (Chwialkowski et al., 2014, Theorem 1). The wild bootstrapped statistics can all be computed efficiently as outlined in Section˜1.2.
@@
The KSD test is known to be consistent achieving pointwise power (Chwialkowski et al., 2016, Proposition 3.2, and Liu et al., 2016, Proposition 4.2), that is, for any fixed alternative, the power of the KSD test converges to 1 asymptotically. In Sections˜6.2 and 6.3, we present asymptotic power guarantees in a setting in which alternatives depend on the sample size and shrink towards the null as it increases.
@@
The permutation-based privatisation and robustisation procedures of Kim and Schrab (2023, Algorithm 1) and Schrab and Kim (2025, Algorithm 1) only work in frameworks where testing the null corresponds exactly to testing exchangeability (see Section˜1). As previously mentioned, the two-sample and independence testing frameworks satisfy this property. However, this is not the case for goodness-of-fit testing. Indeed, since we have access to only one sample (and to the model itself), we are not able to permute samples across distributions as in the two-sample case, and permuting the points within one sample has no effect. So, the goodness-of-fit problem cannot be framed as testing for exchangeability. This is the reason why there are no permutation method in this setting, and why the type I error control only holds asymptotically. This also explains why the procedures of Kim and Schrab (2023, Algorithm 1) and Schrab and Kim (2025, Algorithm 1) cannot be used to construct private and robust KSD tests: a task which remains an open problem.
@@
The choice of base kernel and its bandwidth leads to drastically different Stein kernels, the role of the bandwidth is even more crucial in the goodness-of-fit setting as it affects the derivates of the base kernel which appear in the expression of the Stein kernel. Hence, this choice greatly impacts the power of the KSD test. This kernel selection problem can be solved with the aggregation method (Section˜2) or with kernel pooling (Schrab, 2025b, Section 4). We recommend using these two adaptive procedures, leading to the KSDAgg test (Schrab et al., 2022a, Algorithm 1) and the normalised KSDFuse test.
@@ 6.2 Uniform power against alternatives separated in KSD metric
The KSD test is guaranteed to have power at least $1-\beta$ against all alternatives separated as (Section˜A.1)
$\mathrm{KSD}_{k}\penalty 10000\ \gtrsim\penalty 10000\ \sqrt{\frac{\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta)\big\}}{N}}$ (44)
which is of order $N^{-1/2}$. Using any unnormalised mean kernel pooling (Schrab, 2025b, Section 4) leads to (Section˜A.4)
$\underset{k\in{\mathcal{K}}}{\mathrm{mean}}\penalty 10000\ \mathrm{KSD}_{k}\penalty 10000\ \gtrsim\penalty 10000\ \sqrt{\frac{\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta)\big\}}{N}},$ (45)
while using unnormalised fuse/max kernel pooling leads to (Section˜A.4)
$\max_{k\in{\mathcal{K}}}\,\mathrm{KSD}_{k}\penalty 10000\ \gtrsim\penalty 10000\ \sqrt{\frac{\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta),\,\log(|{\mathcal{K}}|)\big\}}{N}},$ (46)
which, for a collection of kernels of size $|{\mathcal{K}}|=\log N$ (e.g., Schrab et al., 2022a, Theorem 3.5.ii), corresponds to a rate of order $(N/\log\log N)^{-1/2}$. Here, and throughout this section, the fusing parameter $\nu$ is assumed to be greater than $N$ and $\log(|{\mathcal{K}}|)$. In the following, we present results for fuse/max kernel pooling, while mean pooling results can be derived in a similar way to Equation˜45 achieving the same rate as for fixed kernel.
@@
The efficient goodness-of-fit test, based on the block KSD B-statistic (Schrab, 2025b, Equation 70) consisting of $B$ blocks, with fixed kernel $k$, controls the type II error by $\beta$ provided that (Section˜A.5)
$\mathrm{KSD}_{k}\penalty 10000\ \gtrsim\penalty 10000\ \sqrt{\frac{B\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta)\big\}}{N}}.$ (47)
If only one block is used (i.e., $B=1$ corresponding to the complete U-statistic), then the same rate $N^{-1/2}$ as in Equation˜44 is achieved. As the number of blocks $B$ increases, the rate $(B/N)^{1/2}$ gets slower until it no longer converges to zero for $B\asymp N$. The unnormalised fuse/max pooled block KSD statistic over a collection $\mathcal{K}$} of kernels (often of size $|{\mathcal{K}}|=\log(|{\mathcal{D}}|/N)\approx\log(N/B)$ as in Schrab et al. (2022b, Theorem 2.ii) with $|{\mathcal{D}}|=B\lfloor N/B\rfloor^{2}\asymp N^{2}/B$) achieves the same uniform separation rate with an additional $\sqrt{\log|{\mathcal{K}}|}$ term, that is (Section˜A.4)
$\max_{k\in{\mathcal{K}}}\,\mathrm{KSD}_{k}\penalty 10000\ \gtrsim\penalty 10000\ \sqrt{\frac{B\max\!\big\{\!\log(1/\alpha),\,\log(1/\beta),\,\log(|{\mathcal{K}}|)\big\}}{N}}.$ (48)